OpenStudy (anonymous):

Write the slope-intercept equation for the line with slope 1/2 and y-intercept (0, -5) and graph the line.

OpenStudy (anonymous):

help

OpenStudy (nali):

y= mx+b subtitute and then it is easy to graph -5= (1/2) (0) + b

OpenStudy (nali):

b= -5

OpenStudy (nali):

so the equation is y= (1/2)x - 5

OpenStudy (nali):

then just graph it

OpenStudy (anonymous):

thanks

OpenStudy (nali):

u r welcome